Liz needs to buy soft drinks for a party for 30 people. When she gets to the store, she finds that she can buy six cans for $2.10. She assumes that each person will drink two cans. When she checks her wallet, she finds that she has only $20. Does she have enough money to buy the soft drinks? Explain your reasoning.
Found 2 solutions by jim_thompson5910, mananth:
Answer by jim_thompson5910(35256)

If 30 people drink 2 cans each, then she needs 60 cans.


Since "she can buy six cans for $2.10", and she needs 60 cans, she's going to need to buy 10 cases (since 6*10 = 60) which will cost 10*2.1 = 21 dollars total.


Since $21 is more than $20, she does NOT have enough.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim

Answer by mananth(16949)

there are 30 people
two per person
= 60 drinks required.
...
6 drinks cost $2.10
60 dinks will cost
60/6 * 2.10
=10*2.10
= 21
$21
..
She has only $20. Can she buy?
...
m.ananth@hotmail.ca